During 2020, ABC Co. had net income of $900,000. The company has 20,000 shares of 5%, $100 par, convertible preferred stock outstanding, and 1,500, 10% convertible bonds outstanding, which were issued at par. The preferred shares and the convertible bonds were issued in 2017. Each share of convertible preferred stock is convertible into 5 shares of common stock. Each bond is convertible into 90 shares of common stock. The company has a tax rate of 21% and a weighted average shares outstanding of 800,000 throughout 2020.

Instructions:

  1. Compute Basic EPS for 2020
  2. Compute Diluted EPS for 2020

Please type your answers in the text-box below. Please show your work using the EPS formula equations.
Answer #1
It is considered that the Net Income is net of preferred dividend ie. after preferred dividend
Thus, the provided net income is income attributable to the common stock.
Basic EPS (2020) = Net Income / Average shares outstanding
Basic EPS (2020) = 900000 / 800000 = $1.13 per share
Diluted EPS (2020) = (Net Income + Interest cost of bonds + Preferred dividend) / (outstanding shares + convertible shares)
Diluted EPS (2020) =[ 900000 + (1500*1000*10%) + (20000*100*5%)] / [800000 + (20000*5) + (1500*90)]
Diluted EPS (2020) = (900000 + 150000 + 100000) / (800000 + 100000 + 135000) = $1.11 per share